Particular integral of d²y/dx² +3dy/xdx +2/x² = 0 is
1
Expert's answer
2020-08-27T17:00:01-0400

Let:

"y'=v(x)"

then:

"y''=v'"

Let:

"\mu(x)=e^{\int3/xdx}=x^3"

Multiply all terms of the given equation by "\mu(x)" :

"x^3v'+3x^2v=-2x"

Substitute "3x^2=\frac{d}{dx}(x^3)" :

"x^3\frac{dv(x)}{dx}+\frac{d}{dx}(x^3)v(x)=-2x"

Then:

"\frac{d}{dx}(x^3v)=-2x"

"\int\frac{d}{dx}(x^3v)dx=-\int2xdx"

"x^3v=-x^2+c_1"

"v=\frac{dy}{dx}=\frac{-x^2+c_1}{x^3}"

Answer:

"y(x)=\int\frac{-x^2+c_1}{x^3}dx=-lnx-\frac{c_1}{2x^2}+c_2"